OpenStudy (anonymous):

Solve the following system of equations graphically. Click on the graph until the correct solution of the system appears. y = -3 x - y = 8

OpenStudy (mathstudent55):

You need to graph both equations. The first one is a horizontal line that goes through y = -3. The second one can be graphed by finding two points and connecting them with a line. To find two points, let x equal values, such as 0 and 4. Then solve for y. That gives you two points. Plot the points and conect them with a line. The solution of the system of equations is the point where the two lines intersect. Read its coordinates off the graph.

OpenStudy (anonymous):

thank you so much

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

x-y=8 when y=-3,x+3=8,x=8-3=5 when y=0,x=8 AB &CD intersect at C Hence solution is (0,-3)

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

yw